OpenStudy (anonymous):

A restaurant has x floors. The number of tables on each floor is 4 less than the number of floors. The number of chairs with each table is 1 less than the number of tables on each floor. The expression below shows the total number of chairs in the restaurant: x(x - 4)(x - 5) Part A: What does x(x - 4) represent? Part B: What type of polynomial is x(x - 4)(x - 5)? Part C: What is the degree of the polynomial x(x - 4)?

OpenStudy (anonymous):

What do you think about Part A?

OpenStudy (anonymous):

I really dont know

OpenStudy (anonymous):

Im stuck in that one

OpenStudy (jadeishere):

x = floors x - 4 = tables x - 5 = amount of chairs x(x-4) = the amount of tables all together x(x-5) amount of chairs Want to know why?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

they are telling you that the number of tables on each floor is 4 less than the number of floors (x - 4) So, x(x - 4) would be the number of tables in the floor

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Thanks, for part B i got x3−9x2+20x

OpenStudy (anonymous):

So it will be a trinomial right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

trinomial = degree 3

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

Thank u so much

OpenStudy (anonymous):

what about x(x - 4)?

OpenStudy (anonymous):

Wait Part a : The number of tables Part B: Trinomial Part c: 3rd Degree

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

part a you have to mention that are the number of tables on the floor

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Part A: The number of tables on the floor*

OpenStudy (anonymous):

Part B: Trinomial

OpenStudy (anonymous):

Part C: 3rd Degree

OpenStudy (anonymous):

part C is binomial, so degree 2

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

x(x - 4) = x^2 - 4x

OpenStudy (anonymous):

So Part C: 2nd Degree

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank u so much!!!

OpenStudy (anonymous):

no problem